OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (nincompoop):

try

OpenStudy (unklerhaukus):

\[\vec u\cdot\vec v=|u||v|\cos\theta\]